OpenStudy (anonymous):

Given \cos(\alpha)=\frac{\sqrt{17}}{9} for 0<\alpha<\pi/2. Also given \cos(\beta)=\frac{\sqrt{7}}{4} and 0<\beta<\pi/2. Use sum and difference formulas to find the following: Note: You are not allowed to use decimals in your answer. \tan(\alpha + \beta)=

OpenStudy (anonymous):

can you retype your answer? its so hard to understand

OpenStudy (anonymous):

retype the question i mean

OpenStudy (tyteen4a03):

\(\cos(\alpha)=\frac{\sqrt{17}}{9}\) for \(0<\alpha<\pi/2\). Also given \(\cos(\beta)=\frac{\sqrt{7}}{4}\) and \(0<\beta<\pi/2\). \(\tan(\alpha + \beta)=\) Retyped the equation bit.

OpenStudy (anonymous):

\[\cos (\alpha)=\frac{ \sqrt{17} }{ 9 } for 0<\alpha<\frac{ \pi }{ 2}, also given \cos (\beta)=\frac{ \sqrt{7} }{ 4} and 0<\beta<\frac{ \pi }{ 2 }\]

OpenStudy (anonymous):

\[\cos(\alpha+\beta)=\cos(\alpha)\cos(\beta)-\sin(\alpha)\sin(\beta)\]\[=\cos(\alpha)\cos(\beta)-\sqrt{1-\cos^2(\alpha)}\sqrt{1-\cos^2(\beta)}\]Use that to get cos(a+b) \[\tan(\alpha+\beta)=\frac{ \sin(\alpha+\beta) }{ \cos(\alpha+\beta) }=\frac{\sqrt{1-\cos^2(\alpha+\beta)}}{\cos(\alpha+\beta)}\] You work out the details! :)

OpenStudy (anonymous):

ok. thank you